# A cutting-surface consensus approach for distributed robust optimization of multi-agent systems

### v2

@@ 5.2 Analysis of Local Feasibility of The Solutions
Figure 7: Comparison of Algorithm 1 with the cutting-plane consensus algorithm in [30, 29] in terms of solution feasibility of the agent 1. The blue curve indicates the profile of $g_{1}(\hat{z}_{1}^{k},y)$ regarding the uncertainty parameter y, where $\hat{z}_{1}^{k}$ is the resulting solution of the cutting-plane consensus algorithm. Similarly, the black curve is the profile obtained according to the cutting-surface consensus approach.
Table 3: Feasibility of the solution of an agent to its own constraint
@@
In this subsection, we compare Algorithm 1 with the cutting-plane consensus algorithms in [30, 29] to present the superiority of our algorithm in terms of feasibility assurance of the resulting solutions. The cutting-plane consensus algorithms are based on successively populating cutting-planes into the finite set of constraints to externally approximate the (${\rm DRCP}$), which fails to ensure the feasibility of the resulting solutions. By comparison, Algorithm 1 is based on iteratively approximating the (${\rm DRCP}$) by successively reducing the restriction parameters $\epsilon^{k}_{i}$ of the right-hand constraints and expanding the finite sets $Y_{i}^{k}$, which enables each agent to obtain a feasible solution within a finite number of iterations. This result is illustrated by Figure 7 and Table 3.
Additionally, compared with the cutting-plane consensus algorithms in [30, 29], our approach possesses the following three characteristics : 1) Our approach can handle optimization problems with different local cost functions; 2) Our approach eliminates the need to exchange local constraints with neighboring agents, which reduces communication costs and protects the privacy of each agent; 3) The constraints updated by our approach at each iteration are non-polyhedral, which allows us to deal with optimization problems with high-dimensional decision variables more efficiently [31].
@@ 5.3 Computational Performance of Adjusting $\epsilon^{0}_{i}$ and $r$
Figure 8: Effect of turning parameter $\epsilon^{0}_{i}$. The reduction parameter is $r\!=\!2$.
Figure 9: Effect of turning parameter r. The initial restriction parameter is $\epsilon^{0}_{i}\!=\!100$ for any $i\in\mathcal{V}$.
For a fixed reduction parameter $r=2$, the effect of turning parameter $\epsilon^{0}_{i}$ is shown in Figure 8, where $10^{-4}<\epsilon^{0}_{i}<10^{5}$. When the parameter $\epsilon^{0}_{i}$ is small ($\epsilon^{0}_{i}\leq 1$), the number of solvability cuts remains zero as the (${\rm ADRCP^{k}}$) is solvable throughout the iterations of Algrithm 1. Besides, the number of optimality cuts sees a steady increase as $\epsilon^{0}_{i}$ rises in this stage. Conversely, when the parameter $\epsilon^{0}_{i}$ is large ($\epsilon^{0}_{i}>1$), the (${\rm ADRCP^{k}}$) becomes unsolvable in the initial iterations of Algorithm 1, leading to a growth in the number of solvability cuts with $\epsilon^{0}_{i}$. However, the number of optimality cuts remains basically unchanged in this stage. In addition, the number of feasibility cuts is basically unchanged regardless of the value of $\epsilon^{0}_{i}$. To sum up, as $\epsilon^{0}_{i}$ rises, the number of iterations of Algorithm 1 increases due to the increase in optimality cuts (for small $\epsilon^{0}_{i}$) and solvability cuts (for large $\epsilon^{0}_{i}$).
For a fixed restriction parameter $\epsilon^{0}_{i}=100$, the effect of turning parameter $r$ is shown in Figure 9, where $1<r\leq 1000$. When $r$ is small ($r<10$), increasing $r$ leads to a sharp decrease in the quantities of optimality cuts and solvability cuts. Conversely, for larger values of $r$ ($r>10$), the number of optimality cuts and solvability cuts decreases gradually and eventually stabilizes. Regardless of the value of $r$, the number of feasibility cuts remains virtually unchanged. In summary, as $r$ increases, the number of iterations of Algorithm 1 decreases along with the reduction in the quantities of optimality cuts and solvability cuts.
@@
It is worth noting that there is a limit to the computational accuracy of MATLAB. Therefore, we cannot set $\epsilon^{0}_{i}$ too small ($\epsilon^{0}_{i}<10^{-7}$), or $r$ too large ($r>10^{3}$), as this would lead to the number of feasibility cuts and iterations tends to infinity, thus preventing Algorithm 1 from terminating within a finite number of iterations.
@@ 6 Conclusion
A cutting-surface consensus approach is proposed for solving the distributed robust convex optimization problems of multi-agent systems with guaranteed both feasibility and convergence, which is applicable to time-varying unbalanced directed graphs under the assumption of uniformly jointly strong connectivity. Then, it is proved that this approach converges to a feasible solution satisfying consensus and zeroth-order stopping conditions to certain tolerances within a finite number of iterations. Moreover, we use a numerical example to illustrate the effectiveness of the approach.
Direct extensions may lie in the following two aspects. First, in the current cutting-surface consensus approach, we use the distributed projected gradient (DPG) algorithm to solve the approximate problem (${\rm ADRCP^{k}}$). We could also design other constrained distributed optimization algorithms to solve the problem, ensuring that the resulting solutions of these algorithms are feasible to the (${\rm ADRCP^{k}}$). In addition, this paper mainly studies robust optimization with bounded uncertainty, and future research may consider studying more general distributionally robust optimization (DRO) problems with moment uncertainty in multi-agent systems.
@@
In this Appendix, referring to the proof idea of the distributed projected subgradient algorithm under fixed unbalanced graphs [41], we prove the convergence and convergence rate of the proposed DPG algorithm in Section 3.1.
@@ 6.1 Proof of Theorem 1
Let $\left\{\mathcal{G}(t)\right\}$ and $\left\{A(t)\right\}$ satisfy Assumptions 2, 4. For any $s\geq t$, define $A(s:t)=A(s-1)\cdot\cdot\cdot A(t)$ where $A(t:t)=I$ and $a_{ij}(s:t)$ denote the entries of $A(s:t)$. For any $t\geq 0$, there exists a normalized vector $\pi(t)$, i.e., $11_{m}^{\top}\pi(t)=$ 1, such that
@@
    there exist $B>0$ and $10<\lambda<$ 1 such that $|a_{ij}(s:t)-\pi_{j}(t)|\leq B\lambda^{s-t}$ for any $i,j\in\mathcal{V}$ and $s\geq t$.
@@
Let $U$ be the convex hull of $\cup_{i\in\mathcal{V}}\Omega_{i}$, i.e. $U={\rm conv}\big{(}\bigcup_{i\in\mathcal{V}}\Omega_{i}\big{)}$. Suppose that Assumption 3 holds and $\Omega_{i}$ is a compact set, then there exists a constant $R\geq 1$ such that
@@
Let Assumptions 2, 3, 4 be satisfied. Then for the algorithm (3), the following holds for any $v\in\Omega$ and $t\geq 0$:
@@
${\|\tilde{\theta}_{i}(t)-v\|}^{2}\leq\sum_{j=1}^{m}a_{ij}(t){\|(\theta_{j}(t)-v)\|}^{2}.$ (31)
@@
$c^{\top}(v-\tilde{\theta}_{i}(t))$ $=c^{\top}[v-\overline{\theta}(t)+\overline{\theta}(t)-\tilde{\theta}_{i}(t)]$ (32)
$\leq c^{\top}(v-\overline{\theta}(t))+\big{|}c^{\top}\tilde{\theta}_{i}(t)-c^{\top}\overline{\theta}(t)\big{|}$
$\leq\!c^{\top}\!(v\!-\!\overline{\theta}(t))\!+\!C\!\sum_{j=1}^{m}\!a_{ij}(t)\|\theta_{j}(t)-\overline{\theta}(t)\|,$
[…]
$\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}$ (33)
$\leq\sum_{i=1}^{m}\pi_{i}(t+1)\sum_{j=1}^{m}a_{ij}(t){\|\theta_{j}(t)-v\|}^{2}$
$\ \ -\sum_{i=1}^{m}\pi_{i}(t+1)2\alpha(t)c^{\top}(\overline{\theta}(t)-v)$
$\ \ +\sum_{i=1}^{m}\pi_{i}(t+1)2\alpha(t)C\sum_{j=1}^{m}a_{ij}(t)\|\theta_{j}(t)-\overline{\theta}(t)\|$
$\ \ +\sum_{i=1}^{m}\pi_{i}(t+1)\alpha^{2}(t)C^{2}-\sum_{i=1}^{m}\pi_{i}(t+1){\|\phi_{i}(t)\|}^{2}.$
@@
Let Assumptions 2, 3, 4 be satisfied. Then for the algorithm (3), the following relations hold.
@@
Proof 6.24.
Parts (a) and (c) are straightforward. Here, we only prove that part (b) holds. Firstly, we write (3) as $\theta_{i}(t+1)=\sum_{j=1}^{m}a_{ij}(t)\theta_{j}(t)+\epsilon_{i}(t)$, where $\epsilon_{i}(t)\in\mathbb{R}^{n}$ represents an error term. Hence, it holds that $\theta_{i}(t)=\sum_{j=1}^{m}[A(t:0)]_{ij}\theta_{j}(0)+\sum_{r=0}^{t-1}\sum_{j=1}^{m}[A(t:r+1)]_{ij}\epsilon_{j}(r)$. Due to $\overline{\theta}(t)=\sum_{j=1}^{m}\pi_{j}(t)\theta_{j}(t)$ and Lemma 6.18(c), we obtain $\overline{\theta}(t)=\sum_{j=1}^{m}\pi_{j}(0)\theta_{j}(0)+\sum_{r=0}^{t-1}\sum_{j=1}^{m}\pi_{j}(r+1)\epsilon_{j}(r)$. Therefore, $\|\theta_{i}(t)-\overline{\theta}(t)\|\!=\!\|\sum_{j=1}^{m}\big{(}[A(t:0)]_{ij}\!-\!\pi_{j}(0)\big{)}\theta_{j}(0)\!+\!\sum_{r=0}^{t-1}\sum_{j=1}^{m}\big{(}[A(t:r+1)]_{ij}-\pi_{j}(r+1)\big{)}\epsilon_{j}(r)\|\!\leq\!\sum_{j=1}^{m}\big{|}[A(t:0)]_{ij}-\pi_{j}(0)\big{|}\|\theta_{j}(0)\|+\sum_{r=0}^{t-1}\sum_{j=1}^{m}\big{|}[A(t:r+1)]_{ij}-\pi_{j}(r+1)\big{|}\|\epsilon_{j}(r)\|$. By Lemma 6.18(a), we have
@@
Proof 6.26.
It is straightforward to derive from Lemma 7 in [35].
Based on Propositions 6.21 and 6.23. we obtain the following important result, which is crucial for proving the convergence of the distributed projected gradient algorithm.
@@
Let Assumptions 2, 3, 4 be satisfied. If $\left\{\alpha(t)\right\}$ is nonincreasing, then for the algorithm (3), we have
$\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}\!+\!ab\psi(t+1)$ (35)
$\leq\sum_{i=1}^{m}\!\pi_{i}(t){\|\theta_{i}(t)\!-\!v\|}^{2}\!+\!ab\psi(t)\!-\!2\alpha(t)c^{\top}(w(t)\!-\!v)\!+\!D_{5}\alpha^{2}(t)$
$-\!D_{4}\sum_{i=1}^{m}\|\phi_{i}(t)\|^{2}\!+\!D_{13}\alpha(t)\lambda^{t}+D_{23}\alpha(t)\sum_{r=0}^{t-1}\lambda^{t-r-1}\alpha(r)$
@@
$\|w(t)-\overline{\theta}(t)\|\!=\!{\rm dist}(\overline{\theta}(t),\Omega)\!\leq\!R\max_{i\in\mathcal{V}}{\rm dist}(\overline{\theta}(t),\Omega_{i})$ (36)
$\ \leq\!\sum_{i=1}^{m}\frac{R\pi_{i}(t)}{\eta}{\rm dist}(\overline{\theta}(t),\Omega_{i})\!\leq\!\sum_{i=1}^{m}\frac{R\pi_{i}(t)}{\eta}\|\theta_{i}(t)\!-\!\overline{\theta}(t)\|,$
where the second inequality follows from $\pi_{i}(t)/\eta\geq 1$ (cf. Lemma 6.18(b)), and the last inequality is satisfied as ${\rm dist}(\overline{\theta}(t),\Omega_{i})\leq\|\theta_{i}(t)-\overline{\theta}(t)\|$ (cf. formula (2) in [35]). Thus, $c^{\top}(v-\overline{\theta}(t))\leq c^{\top}(v-w(t))+\frac{RC}{\eta}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|$. Using this bound for (28), we obtain
$\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}$ (37)
$\leq\sum_{i=1}^{m}\pi_{i}(t){\|\theta_{i}(t)-v\|}^{2}-2\alpha(t)c^{\top}(w(t)-v)+C^{2}\alpha^{2}(t)$
$+D_{3}\alpha(t)\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|-\sum_{i=1}^{m}\pi_{i}(t+1)\|\phi_{i}(t)\|^{2},$
@@
$\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}+ab\psi(t+1)$ (38)
$\leq\sum_{i=1}^{m}\pi_{i}(t){\|\theta_{i}(t)\!-\!v\|}^{2}\!+\!ab\psi(t)\!+\!ab(\lambda\!-\!1)\psi(t)\!+\!ab\alpha(t)\beta(t)$
$-\!2\alpha(t)c^{\top}(w(t)\!-\!v)\!+\!C^{2}\alpha^{2}(t)-\!\sum_{i=1}^{m}\pi_{i}(t+1)\|\phi_{i}(t)\|^{2}\!$
$+\!D_{13}\alpha(t)\lambda^{t}\!+\!D_{23}\alpha({t})\sum_{r=0}^{t-1}\lambda^{t-r-1}\alpha(r)\!+\!D_{3}B\psi(t).$
@@
It now remains to demonstrate that $\delta=0$. By using the triangle and Cauchy-Schwarz inequalities, it can be confirmed that $\|\theta_{i}(t)-\theta^{*}\|^{2}\leq 3(\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}+\|\overline{\theta}(t)-w(t)\|^{2}+\|w(t)-\theta^{*}\|^{2})$. Next, since $\|\overline{\theta}(t)-w(t)\|\leq\frac{R}{\eta}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|$ (cf. inequality (36)), we have $\|\overline{\theta}(t)-w(t)\|^{2}\leq\frac{R^{2}}{\eta^{2}}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}$ by the convexity of $\|\cdot\|^{2}$. As a result, $\frac{\|\theta_{i}(t)-\theta^{*}\|^{2}}{3}\leq\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}+\frac{R^{2}}{\eta^{2}}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}+\|w(t)-\theta^{*}\|^{2}$. Multiplying both sides by $\pi_{i}(t)$ then summing over $i\in\mathcal{V}$ yields the follwing, with $R^{\prime}=1+R^{2}/\eta^{2}$:

### v3

@@ 4.2 Analysis of Local Feasibility of The Solutions
Figure 6: Comparison of cutting-surface consensus algorithm (i.e., Algorithm 2) with the cutting-plane consensus algorithm in [27, 26] in terms of solution feasibility of the agent 1. The blue curve indicates the profile of $g_{1}(\hat{z}_{1}^{k},y_{1})$ regarding the uncertainty parameter $y_{1}$, where $\hat{z}_{1}^{k}$ is the resulting solution of the cutting-plane consensus algorithm. Similarly, the black curve is the profile obtained according to the cutting-surface consensus approach.
Table 3: Local Feasibility of the solution
@@
In this subsection, we compare Algorithm 2 with the cutting-plane consensus algorithms in [27, 26] to present the superiority of our algorithm in terms of feasibility assurance of the resulting solutions. The cutting-plane consensus algorithms are based on successively adding linear cutting-planes into the finite set of constraints to externally approximate the ($\rm DRCP$}), which fails to ensure the feasibility of the resulting solutions. By comparison, Algorithm 2 is based on iteratively approximating the ($\rm DRCP$}) by successively reducing the restriction parameters $\epsilon^{k}_{i}$ of the right-hand constraints and adding convex, possibly nonlinear cutting-surfaces into the existing finite sets of constraints, which enables each agent to obtain a local feasible solution within a finite number of iterations. This result is illustrated by Figure 6 and Table 3.
@@ 4.3 Computational Performance of Adjusting $\epsilon^{0}_{i}$ and $r$
Figure 7: Effect of tuning parameter $\epsilon^{0}_{i}$. The reduction parameter is r=2r\!=\!2.
Figure 8: Effect of tuning parameter $r$. The initial restriction parameter is $\epsilon^{0}_{i}\!=\!100$ for any $i\in\mathcal{V}$.
For a fixed reduction parameter $r=2$, the effect of tuning parameter $\epsilon^{0}_{i}$ is shown in Figure 7, where $10^{-4}<\epsilon^{0}_{i}<10^{5}$. When the parameter $\epsilon^{0}_{i}$ is small ($\epsilon^{0}_{i}\leq 1$), the number of Case I remains zero as the ($\rm ADRCP^{k}$}) is feasible throughout the iterations of Algrithm 2. In addition, the number of Case III sees a steady increase as $\epsilon^{0}_{i}$ rises in this stage. Conversely, when the parameter $\epsilon^{0}_{i}$ is large ($\epsilon^{0}_{i}>1$), the ($\rm ADRCP^{k}$}) becomes unfeasible in the initial iterations of Algorithm 2, leading to a growth in the number of Case I with $\epsilon^{0}_{i}$. However, the number of Case III remains basically unchanged in this stage. Moreover, the number of Case II is basically unchanged regardless of the value of $\epsilon^{0}_{i}$. To sum up, as $\epsilon^{0}_{i}$ rises, the number of iterations of Algorithm 2 increases due to the increase in Case III (for small $\epsilon^{0}_{i}$) and Case I (for large $\epsilon^{0}_{i}$).
For a fixed restriction parameter $\epsilon^{0}_{i}=100$, the effect of tuning parameter $r$ is shown in Figure 8, where $1<r\leq 10^{3}$. When $r$ is small ($r<10$), increasing $r$ leads to a sharp decrease in the quantities of Case III and Case I. Conversely, for larger values of $r$ ($r>10$), the number of Case III and Case I decreases gradually and eventually stabilizes. Regardless of the value of $r$, the number of Case II remains virtually unchanged. In summary, as $r$ increases, the number of iterations of Algorithm 2 decreases along with the reduction in the quantities of Case III and Case I.
@@
There is a limit to the computational accuracy of MATLAB. Therefore, we cannot set $\epsilon^{0}_{i}$ too small ($\epsilon^{0}_{i}<10^{-7}$), or $r$ too large ($r>10^{3}$), as this would lead to an infinite number of Case II and iterations, thus preventing Algorithm 2 from terminating within a finite number of iterations.
@@ 5 Conclusion
A distributed cutting-surface consensus approach is proposed for solving the distributed robust convex optimization problems of multi-agent systems with guaranteed local feasibility and finite-time termination, which is applicable to time-varying unbalanced directed graphs under the assumption of uniformly jointly strong connectivity. Then, it is proved that this approach enables each agent to locate a local feasible solution satisfying consensus and zeroth-order stopping conditions to certain tolerances within a finite number of iterations. Moreover, we use a numerical example to illustrate the effectiveness of the approach.
Direct extensions may lie in the following four aspects. Firstly, in the current cutting-surface consensus approach, we use the distributed projected gradient (DPG) algorithm to solve the approximate problem ($\rm ADRCP^{k}$}), where the convergence rate of the DPG algorithm is sublinear. Future research could consider designing optimization algorithms with faster convergence rates (such as linear or exponential convergence) to replace the DPG algorithm, thereby improving the efficiency of the cutting-surface consensus algorithm. Secondly, the stopping criterion of this paper is not sufficient for finite-time convergence, it is sufficient for finite-time termination and enabling each agent to obtain an approximate optimal solution that satisfies consensus and zeroth-order stopping conditions to certain tolerances. Ongoing work will consider propose a novel distributed robust convex optimization algorithm that finite-time converges to a feasible solution with guaranteed sub-optimality level. In addition, it would be interesting to conduct a thorough theoretical and computational analysis of convergence rate of the distributed cutting-surface consensus algorithm. Finally, this paper mainly considers the (DRCP) with only one semi-infinite constraint for each agent. Future work will consider a more complex case that (DRCP) includes multiple semi-infinite constraints for each agent.
Acknowledgments
We are grateful to the associate editor and the anonymous reviewers whose comments lead to substantial improvement of this article, especially Reviewers 2 and 3.
@@ Acknowledgments
In this Appendix, referring to the proof idea of [33], we prove the convergence and convergence rate of the DPG algorithm proposed in Section 3.1.
@@ 5.1 Proof of Theorem 1
Let $\left\{\mathcal{G}(t)\right\}$ and $\left\{A(t)\right\}$ satisfy Assumptions 2, 4. For any $s\geq t$, define $A(s:t)=\prod_{k=t}^{s-1}A(k)$ where $A(t:t)=I$ and $a_{ij}(s:t)$ denotes the entries of $A(s:t)$. For any $t\geq 0$, there exists a normalized vector $\pi(t)$, i.e., $1_{m}^{\top}\pi(t)=1$, such that
@@
    there exist $B>0$ and $0<\lambda<1$ such that $|a_{ij}(s:t)-\pi_{j}(t)|\leq B\lambda^{s-t}$ for any $i,j\in\mathcal{V}$ and $s\geq t$.
@@
Let $U$ be the convex hull of $\cup_{i\in\mathcal{V}}\Omega_{i}$, i.e. $U={\rm conv}(\cup_{i\in\mathcal{V}}\Omega_{i})$. Suppose that Assumption 3 holds and $\Omega_{i}$ is a compact set, then there exists a constant $R\geq 1$ such that
@@
Let Assumptions 24 be satisfied. Then for the algorithm (3), the following holds for any $v\in\Omega$ and $t\geq 0$:
@@
${\|\tilde{\theta}_{i}(t)-v\|}^{2}\leq\sum_{j=1}^{m}a_{ij}(t){\|\theta_{j}(t)-v\|}^{2}.$ (29)
@@
$\hskip-19.91692pt\begin{aligned} &&c^{\top}(v-\tilde{\theta}_{i}(t))&\leq c^{\top}(v-\overline{\theta}(t))+\big|c^{\top}\tilde{\theta}_{i}(t)-c^{\top}\overline{\theta}(t)\big|\\ &&&\leq\!c^{\top}\!(v\!-\!\overline{\theta}(t))\!+\!C\!\sum_{j=1}^{m}\!a_{ij}(t)\|\theta_{j}(t)-\overline{\theta}(t)\|,\end{aligned}$ (30)
where the first inequality derives from the triangle inequality, and the last inequality follows from the Cauchy-Schwarz inequality and the convexity of $\|\cdot\|$. Using the relations of (29) and (30), we can obtain the upper bound of (28). Then, we take this upper bound into (27) and execute a weighted sum on both sides of the (27), yielding:
$\hskip-14.22636pt\begin{aligned} &&&\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}\\ &&&\leq\sum_{i=1}^{m}\pi_{i}(t+1)\sum_{j=1}^{m}a_{ij}(t){\|\theta_{j}(t)-v\|}^{2}\\ &&&-\sum_{i=1}^{m}\pi_{i}(t+1)2\alpha(t)c^{\top}(\overline{\theta}(t)-v)\\ &&&+\sum_{i=1}^{m}\pi_{i}(t+1)2\alpha(t)C\sum_{j=1}^{m}a_{ij}(t)\|\theta_{j}(t)-\overline{\theta}(t)\|\\ &&&+\sum_{i=1}^{m}\pi_{i}(t+1)\alpha^{2}(t)C^{2}-\sum_{i=1}^{m}\pi_{i}(t+1){\|\phi_{i}(t)\|}^{2}.\end{aligned}$ (31)
@@
Let Assumptions 24 be satisfied. Then for the algorithm (3), the following relations hold.
@@
Proof 5.23.
Parts (a) and (c) are straightforward. Here, we only prove that part (b) holds. Firstly, we rewrite (3) as $\theta_{i}(t+1)=\sum_{j=1}^{m}a_{ij}(t)\theta_{j}(t)+\epsilon_{i}(t)$, where $\epsilon_{i}(t)\in\mathbb{R}^{n}$ represents an error term. Hence, it holds that $\theta_{i}(t)=\sum_{j=1}^{m}[A(t:0)]_{ij}\theta_{j}(0)+\sum_{r=0}^{t-1}\sum_{j=1}^{m}[A(t:r+1)]_{ij}\epsilon_{j}(r).$ Due to $\overline{\theta}(t)=\sum_{j=1}^{m}\pi_{j}(t)\theta_{j}(t)$ and Lemma 5.17(c), we obtain $\overline{\theta}(t)=\sum_{j=1}^{m}\pi_{j}(0)\theta_{j}(0)+\sum_{r=0}^{t-1}\sum_{j=1}^{m}\pi_{j}(r+1)\epsilon_{j}(r).$ Therefore,
$\|\theta_{i}(t)-\overline{\theta}(t)\|$ $=\bigg\|\sum_{j=1}^{m}\big([A(t:0)]_{ij}-\pi_{j}(0)\big)\theta_{j}(0)$
$+\sum_{r=0}^{t-1}\sum_{j=1}^{m}\big([A(t:r+1)]_{ij}-\pi_{j}(r+1)\big)\epsilon_{j}(r)\bigg\|$
$\leq\sum_{j=1}^{m}\big|[A(t:0)]_{ij}-\pi_{j}(0)\big|\|\theta_{j}(0)\|$
$+\sum_{r=0}^{t-1}\sum_{j=1}^{m}\big|[A(t:r+1)]_{ij}-\pi_{j}(r+1)\big|\|\epsilon_{j}(r)\|.$
By Lemma 5.17(a), we have
@@
Proof 5.25.
It can be derived from Lemma 7 in [31].
Based on Propositions 5.20 and 5.22. we obtain the following important result, which is crucial for proving the convergence of the distributed projected gradient algorithm.
@@
Let Assumptions 24 be satisfied. If $\left\{\alpha(t)\right\}$ is nonincreasing, then for the algorithm (3), we have
$\hskip-5.69054pt\begin{aligned} &\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}\!+\!ab\psi(t+1)\\ &\leq\sum_{i=1}^{m}\!\pi_{i}(t){\|\theta_{i}(t)\!-\!v\|}^{2}\!+\!ab\psi(t)\!-\!2\alpha(t)c^{\top}\!(w(t)\!-\!v)\!+\!D_{5}\alpha^{2}(t)\\ &-\!D_{4}\sum_{i=1}^{m}\|\phi_{i}(t)\|^{2}\!+\!D_{13}\alpha(t)\lambda^{t}+D_{23}\alpha(t)\sum_{r=0}^{t-1}\lambda^{t-r-1}\alpha(r),\end{aligned}$ (33)
@@
$\hskip-11.38109pt\begin{aligned} &\|w(t)-\overline{\theta}(t)\|\!=\!{\rm dist}(\overline{\theta}(t),\Omega)\!\leq\!R\max_{i\in\mathcal{V}}{\rm dist}(\overline{\theta}(t),\Omega_{i})\\ &\ \leq\!\sum_{i=1}^{m}\frac{R\pi_{i}(t)}{\eta}{\rm dist}(\overline{\theta}(t),\Omega_{i})\!\leq\!\sum_{i=1}^{m}\frac{R\pi_{i}(t)}{\eta}\|\theta_{i}(t)\!-\!\overline{\theta}(t)\|,\end{aligned}$ (34)
where the second inequality follows from $\pi_{i}(t)/\eta\geq 1$ (cf. Lemma 5.17(b)), and the last inequality is satisfied as $\theta_{i}(t)\in\Omega_{i}$. Thus, $c^{\top}(v-\overline{\theta}(t))\leq c^{\top}(v-w(t))+\frac{RC}{\eta}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|$. Using this bound for (26), we obtain
$\hskip-11.38109pt\begin{aligned} &\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}\\ &\leq\sum_{i=1}^{m}\pi_{i}(t){\|\theta_{i}(t)\!-\!v\|}^{2}\!-\!2\alpha(t)c^{\top}(w(t)\!-\!v)\!+\!C^{2}\alpha^{2}(t)\\ &+D_{3}\alpha(t)\!\sum_{i=1}^{m}\!\pi_{i}(t)\|\theta_{i}(t)\!-\!\overline{\theta}(t)\|\!-\!\sum_{i=1}^{m}\pi_{i}(t+1)\|\phi_{i}(t)\|^{2},\end{aligned}$ (35)
@@
$\hskip-11.38109pt\begin{aligned} &\sum_{i=1}^{m}\pi_{i}(t+1){\|\theta_{i}(t+1)-v\|}^{2}+ab\psi(t+1)\\ &\leq\sum_{i=1}^{m}\pi_{i}(t){\|\theta_{i}(t)\!-\!v\|}^{2}\!+\!ab\psi(t)\!+\!ab(\lambda\!-\!1)\psi(t)\!+\!ab\alpha(t)\beta(t)\\ &-\!2\alpha(t)c^{\top}(w(t)\!-\!v)\!+\!C^{2}\alpha^{2}(t)-\!\sum_{i=1}^{m}\pi_{i}(t+1)\|\phi_{i}(t)\|^{2}\!\\ &+\!D_{13}\alpha(t)\lambda^{t}\!+\!D_{23}\alpha({t})\sum_{r=0}^{t-1}\lambda^{t-r-1}\alpha(r)\!+\!D_{3}B\psi(t).\end{aligned}$ (36)
@@
It now remains to demonstrate that $\delta=0$. By using the triangle and Cauchy-Schwarz inequalities, it can be confirmed that $\|\theta_{i}(t)-\theta^{*}\|^{2}\leq 3(\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}+\|\overline{\theta}(t)-w(t)\|^{2}+\|w(t)-\theta^{*}\|^{2})$. Next, by the inequality (34), we have $\|\overline{\theta}(t)-w(t)\|^{2}\leq\frac{R^{2}}{\eta^{2}}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}$ by the convexity of $\|\cdot\|^{2}$. As a result, $\frac{\|\theta_{i}(t)-\theta^{*}\|^{2}}{3}\leq\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}+\frac{R^{2}}{\eta^{2}}\sum_{i=1}^{m}\pi_{i}(t)\|\theta_{i}(t)-\overline{\theta}(t)\|^{2}+\|w(t)-\theta^{*}\|^{2}$. Multiplying both sides by $\pi_{i}(t)$ then summing over $i\in\mathcal{V}$ yields the follwing, with $R^{\prime}=1+R^{2}/\eta^{2}$: